If PQ = x – 5, QR = 2x – 1, and PR = 21, find x. x =


1
Expert's answer
2021-09-01T15:03:29-0400

We know PQ+QR=PR

So, "(x-5)+(2x-1)=21"

"So, 3x-6=21"

So "3x=21+6=27"

So "x=\frac {27} {3} =9" (Ans.)